OpenStudy (anonymous):

how to solve (8)3^x = 5 ?

OpenStudy (anonymous):

\[(8)3^{x} = 5\] , find x.

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

sure its , \(8\times 3^x = 5\) ? take log on both sides, what do u get ?

OpenStudy (kira_yamato):

@hartnn Would it be more intuitive if we divide throughout by 8 first?

hartnn (hartnn):

yes, dividing by 8 is better step

OpenStudy (anonymous):

my answer is x = 1.62 , im not sure whether correct or not. @hartnn

hartnn (hartnn):

i don't think thats correct...how you got that value ?

hartnn (hartnn):

and you will get value in terms of log.... unless i misintrepret the question...

OpenStudy (anonymous):

\[\log 8 \ * log 3^{x} = \log 5\] log 8 * x log 3 = log 5 x log 3 = log 5/log 8 x = log 5 / log 8 / log 3 x = 1.62 these are my steps

OpenStudy (anonymous):

@jayjayokocha divide by 8 on both side. take log base 10. put value of log5,log5/8(=log5-log8). do it, get ASNWER.

hartnn (hartnn):

you are not appling log rules correctly.

hartnn (hartnn):

lets first divide by 8 on both sides, 3^x = 5/8 now take log on both sides

hartnn (hartnn):

use, \(\Large \log a^b=b \log a\)

OpenStudy (anonymous):

so, now i got xlog3 = log 5 - log 8 , what's next ?

hartnn (hartnn):

just divide both sides by log 3 and you would have got 'x' thats it!

OpenStudy (anonymous):

answer is -0.428 ?

hartnn (hartnn):

YES. thats the approximate answer :)

OpenStudy (anonymous):

just a quick one , 2^x = 5 answer is x = 2.32 ? @hartnn

hartnn (hartnn):

yes, thats also correct :)